OpenStudy (maheshmeghwal9):

A stone sliding over a horizontal ice surface stops after moving a distance S=48m. Determine the initial velocity of the stone 'u' if the force of sliding friction of the stone against the ice is 0.06 of the weight of the stone.

OpenStudy (maheshmeghwal9):

@FoolAroundMath @Australopithecus Please help:)

OpenStudy (maheshmeghwal9):

Its answer is given u = 7.56 m/s but I don't know how to get it???????

OpenStudy (maheshmeghwal9):

@Vaidehi09 :)

OpenStudy (foolaroundmath):

Can you find the frictional force and correspondingly the acceleration of the block ? Then use \(v^{2} - u^{2} = 2as \) to find \(u\)

OpenStudy (anonymous):

u have v = 0, s = 48m and F = 0.06mg....this force offers the deceleration. so 0.06mg = -ma a = -0.06g sub in v^2 = u^2 + 2as.

OpenStudy (maheshmeghwal9):

how did u gt 0.06mg=-ma ??????

OpenStudy (anonymous):

the only force acting in the line of motion is the friction.

OpenStudy (maheshmeghwal9):

isn't there\[mg \sin \theta\]also acting????????????

OpenStudy (anonymous):

why? its a horizontal surface...not an inclined one.

OpenStudy (maheshmeghwal9):

would u plz make the figure for me?

OpenStudy (foolaroundmath):

frictional force = 0.06*weight = 0.06mg This provides the deceration. So, 0.06mg = -ma and why \(mg\sin{\theta}\)? Its a horizontal surface

OpenStudy (anonymous):

work done by friction is equal to cange is kinetic energy.... as final velocity is 0.... and friction is 0.06mg \[\frac{1}{2}mv^{2}=0.06mg(s)\]

OpenStudy (anonymous):